OpenStudy (anonymous):

can someone explain how to find values of the six trigonometric functions when given: "Θ lies in Quadrant II. tan(Θ)=-24/7"

OpenStudy (anonymous):

what the answer?

OpenStudy (anonymous):

yeah, i'm not sure how to find it and i have four other problems for homework like it so i was wondering if someone could explain. do you knw?

OpenStudy (anonymous):

u want the value of \[\Theta\]??

OpenStudy (anonymous):

no i need to find the other six trigonometric functions by using the function of tan(Θ)

OpenStudy (mertsj):



OpenStudy (mertsj):

\[\sin \theta=\frac{side opposite}{hypotenuse}=\frac{24}{25}\]

OpenStudy (mertsj):

\[\cos \theta=\frac{side adjacent}{hypotenuse}=\frac{-7}{25}\]

OpenStudy (anonymous):

you just use that with all of the functions? i have them written down. (: but how did you get 25?

OpenStudy (mertsj):

\[\tan \theta=\frac{-24}{7}\]

OpenStudy (mertsj):

I got 25 by using the Pythagorean Theorem

OpenStudy (mertsj):

now the other functions are the reciprocals of those already given:

OpenStudy (mertsj):

\[\cot \theta=\frac{1}{\tan \theta}=\frac{-7}{24}\]

OpenStudy (mertsj):

\[\sec \theta=\frac{1}{\cos \theta}=\frac{-25}{7}\]

OpenStudy (mertsj):

\[\csc \theta=\frac{1}{\sin \theta}=\frac{24}{25}\]

OpenStudy (anonymous):

thank you ! i'm still a little confuzed about the 25. \[a ^{2}+b ^{2}=c ^{2}\] ? right?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

OH i'm sorry i get it now. i wasn't square rooting the final answer. thank you so much !

OpenStudy (mertsj):

\[7^2+24^2=c^2=49+576=625\]

OpenStudy (mertsj):

\[c^2=625\]

OpenStudy (mertsj):

\[c=\sqrt{625}=25\]

OpenStudy (anonymous):

thank you!